OpenStudy (cybershadow):

In an oscillating L-C circuit in which C=4.00 micro F, the maximum potential difference across the capacitor during the oscillations is 1.5 V and the maximum current through the inductor is 50.0 mA What is the inductance L? Plz give a detailed solution or atleast a hint to how to solve it.Thanx in advance

OpenStudy (cybershadow):

any1 Here? I m totally confuzzled with this question,the things i know are - 1) \[f = 1/2\pi \sqrt{LC}\] 2) e = Ldi/dt= q/C Cant get a clue of how to find L :| need a hint ,

OpenStudy (anonymous):

potentiel difference across a capacitor=q/C u can find maximum charge on capacitor from this now use the equation-----> Ldi/dt=q/c. i m not that sure but see if this can help u a bit

OpenStudy (cybershadow):

yeah i was thinking of using that equation , but again i dont know the value of di/dt

OpenStudy (anonymous):

use energy formula. the total energy in an L-C circuit is=(1/2)CV^2=(1/2)LI^2 (I=maximum current) from here u can find L

OpenStudy (cybershadow):

Yeah , thats right Approach ,Thanx , almost forgot the Energy concept , Thanks a lot :)

OpenStudy (anonymous):

ur welcome